Why is sometimes component of some forces added in the same direction and some are equated to each other??? Can't understand the concept...please help!!

Grade:11

3 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Anurag

We have to always form a equation i.e. F=ma    If in case a=0 then forces are equated to get unknown forces and if two forces are in same direction then we need to add. If forces are in opposite direction then also we have to which looks like they are equated. e.g. If F1 and F2 are acting in opposite direction and acceleration is zero then F1+(-F2)=0 hence F1=F2.

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

vikas askiitian expert
13 years ago

by vector law of addition net force = (F12+ F22+2F1F2COS@)1/2

when forces are in same direction along the same line then @ becomes 0 & net resultant R

R =F1+F2 ........1

when they act in opposite direction then @ becomes 180 & net resultant R

R =F1-F2 ......2

in this case if magnitude of both forces is same then resultant becomes zero...

eg)force is applied on a block but it dosent move due to friction ,here friction force & applied forces are in opposite direction & equal in magnitude....

Anurag Roy
13 years ago

Thank you vikas sir and akash sir....both of you helped me a lot in the concept, hope, to have such help throughout my preparation.